The balance sheet of a sole trader revealed the following position.
Fixed Assets
Land and buildings
R
9 170
R
Equipment at cost less depreciation 15 570
Goodwill at cost less amounts written off 10 250 34 990
Investment in shares at cost 2 200

Current Assets
R
R
Trading stock
45 840
Sundry debtors less bad debts written off
19 350
Cash at bank
10 920
76 110
113 300
=
A
Less: Current Liabilities
R
Sundry creditors and accruals
77 600
=
L
35 700
Owner's Interest
R
R
Capital
28 400
Profit for the year
7 300
35 700
=
0
Required
Answer the following questions:
(a) Are the assets correctly classified? If not, re-arrange them and give your reasons for doing this.
(b) What is the current ratio? Is it satisfactory?
(c) What is the acid-test ratio? Is it satisfactory?
(d) Is the business solvent? Give your reasons.
(e) Is the business overtrading? Give your reasons.
(f) What is the profitability ratio of the business?
1
Expert's answer
2020-04-01T09:27:30-0400

a)Fixed Assets:

Goodwill at cost less amounts written off

Land and buildings

Equipment at cost less depreciation

Investment in shares at cost

Current Assets:

Sundry debtors less bad debts written off 

Trading stock 

Cash at bank

by degree of liquidity are assets


b) "current ratio=\frac{current assets} {current liabilities}=\frac{76110}{77600}=0.98"


this is unsatisfying

c)"acid-test ratio=\frac{current assets-inventory}{Current Liabilities}=\frac{76110}{77600}=0.98"

this is unsatisfying

d) The value of the above calculated ratios is below 1, this indicates a high financial risk - the company is not able to consistently pay current bills. Also available Sundry debtors less bad debts written off and  significant level Sundry creditors and accruals. 

e) No significant increase in current ratio. But Great level Trading stock and Sundry creditors and accruals, there may be a significant increase in the relationship between days of accounts payable and days of receivables. Can say that business is overtrading.

f) we calculate the following profitability indicators

"ROCE=\frac{Profit for the year}{Capital} =\frac{7300}{28400}=0.26"

"ROTA =\frac{profit for the year}{total assets}=\frac{7300}{34990+76110}=\frac{7300}{111110}=0.066"

"RCA=\frac{profit for the year}{current assets}=\frac{7300}{76110}=0.09"

"RFA=\frac{profit for the year}{fixet assets}=\frac{7300}{34990}=0.21"

Profitability rates are very low.